Calculate the pH of the solution after the addition of each of the given amounts of 0.0637 M HNO30.0637 M HNO3 to a 60.0 mL60.0 mL solution of 0.0750 M0.0750 M aziridine. The p?apKa of aziridinium is 8.04.

What is the pH of the solution after the addition of 74.9" mL HNO3?
Answer #1

60.0 mL solution of 0.0750 M aziridine = 0.060 L * 0.0750 mole / L = 0.0045 mole.

and

74.9" mL of  0.0637 M HNO3 = 0.0749 L *  0.0637 mole / L = 0.00477 mole.

excess HNO3 = (0.00477 - 0.0045) = 2.71 * 10^-4 mole.

total volume = (60.0 + 74.9) = 134.9 ml = 0.1349 L

molarity of excess HNO3 = 2.71 * 10^-4 mole / 0.1349 L = 2.01 * 10^-3 M

pH = - log [H+]

or

PH = - log (2.01 * 10^-3 )

or

PH = 2.70